40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 20 00:59:02 2011, kiwamu, Summary, General, Weekly report 
    Reply  Mon Jun 27 10:04:29 2011, kiwamu, Summary, General, Weekly report 
       Reply  Mon Jul 4 14:27:35 2011, kiwamu, Summary, General, Weekly report 
          Reply  Mon Jul 25 11:02:19 2011, kiwamu, Summary, General, Weekly report 
             Reply  Mon Jul 25 11:04:22 2011, kiwamu, Summary, General, Weekly report 
                Reply  Mon Jul 25 11:06:38 2011, kiwamu, Summary, General, Weekly report 
                   Reply  Mon Aug 1 22:30:59 2011, kiwamu, Summary, General, Weekly report 
Message ID: 4835     Entry time: Mon Jun 20 00:59:02 2011     Reply to this: 4890
Author: kiwamu 
Type: Summary 
Category: General 
Subject: Weekly report 
This is a summary for the week ending June 19th. Feel free to edit this entry.
(Number of elog entries = 27)

* Refinement of LSC screen
    -> Kissel buttons and some indicators were newly installed
    -> A script to autonatically generate kissel buttons was made

* New BIO installed on ETMY

* LightWave for ABSL
    -> taken out from the MOPA box and put on the AP table with temporary use of the Y end laser controller
 
* Shipping 2 RFPDs to LLO
 
* LEDs on the BIO for the vertex suspensions were blown
    -> fixed and re-installed. A test script will be prepared
 
* PEM AA board was fixed
 
* A plot of the MICH noise was produced for the first time
 
* Schnupp asymmetry measurement : Las = 3.64+/-0.32cm
 
* The photo diode on WFS2 has been replaced by YAG-444-4A
 
* SUS binary IO crates were taken out
 
* Fiber died
     ->C1LSC was unable to communicate to C1SUS. Installing a new copper Dolphine fixed the issue.
 
* SURF students came
ELOG V3.1.3-